jim is seven years older than janice. the sum of there wages is 41. how old is each person
Answer by likaaka(51)

You must set up a system of equations. Begin by naming your variables ie.
x is Jim's age &
y is Janice's age
If Jim is 7 yrs older than Janice, your first equation is x = y + 7
If the sum of their ages is 41, then x + y = 41
Here is your system of equations
x = y + 7
x + y = 41
Now substitute the value of x into the second equation to solve for y
(y + 7) + y = 41 **the parenthesis aren't needed here but I used them to illustrate the substitution
Now combine like terms and solve for y
2y + 7 = 41, subtract 7 from both sides
2y + 7 - 7 = 41 - 7, simplify
2y = 34, divide 2 from both sides
(2y)/2 = 34/2
y = 17, so Janice is 17 yrs old
Given x = y + 7 and y = 17, then
x = 17 + 7
x = 24, so Jim is 24 yrs old